# SEDMamba: Enhancing Selective State Space Modelling with Bottleneck Mechanism and Fine-to-Coarse Temporal Fusion for Efficient Error Detection in Robot-Assisted Surgery

### v1

@@ III-B Comparison with the State-of-the-Art
Figure 4: Visualization results of the proposed SEDMamba (green line) and the second-best method Mamba (blue line) on all 8 complete test videos. Each row represents one whole test video. The X-axis is the frame index while the Y-axis is the error probability output by the models. The red translucent background indicates error frames, while the white background indicates normal frames, i.e. ground truth of each frame. Different coloured bars on the X-axis represent different gestures. Red and black boxes are used to highlight significant results.
III-C Computational Efficiency
Given a video spatial embedding sequence $\mathcal{E}\in\mathcal{R}^{L\times D}$, the computational complexity of a single-head self-attention Transformer and a selective SSM in SEDMamba are $O(L^{2}D+LD^{2})$ and $O(LDN)$, respectively. Computationally, self-attention is quadratic to sequence length $L$, while selective SSM is linear to $L$. This computational efficiency makes SEDMamba scalable for gigapixel applications with large video lengths. Besides, as shown in Table II and Table III, SEDMamba achieves the highest performance yet maintains the lowest parameters and FLOPs. Specifically, compared to other selective SSM-based methods, Mamba and Vim, our SEDMamba obtains 58.23% and 75.79% model parameter reductions, and 51.79% and 70.90% FLOPs reductions, respectively. Moverover, the parameters and FLOPs of our SEDMamba are only 27.38% and 9.07% of those of the Transformer-based method ASFormer, respectively, yet it achieves a performance gain of 2.44% in AUC and 8.08% in AP. Compared to methods based on 1D and dilated convolution, i.e. TeCNO, MS-TCN and MS-TCN++, our method gets far more performance improvement with a smaller complexity. This is attributed to the bottleneck construction of BMSS blocks, which compresses the spatial dimension at the input and restores it at the output for computational saving, as well as the FCTF therein BMSS, which provides fine-to-coarse temporal features for selective SSM to model long-term dependencies effectively.
@@ III-C Computational Efficiency
III-D Ablation Study
III-D1 Effectiveness of Key Components in BMSS Block
We study the effect of bottleneck mechanism and FCFT in BMSS blocks. As shown in Table IV, whereby removing each component leads to performance degradation. The bottleneck mechanism yields performance gains of +1.64% in AUC and +3.69% in AP, while model parameters and FLOPs are significantly decreased by 82.98% and 78.83%, which guards against overfitting and enhances model efficiency. Besides, such a bottleneck mechanism compresses features on the spatial dimension, which can mitigate the impact of redundant spatial noise on long-term dependency modelling. Introducing FCTF to bottleneck mechanism leads to a slight increase in computational complexity, which is acceptable given the substantial performance gains of +1.58% in AUC and +3.58% in AP. This underscores the importance of providing multi-scale temporal information for selective SSM in the surgical error detection task.
@@ G Params (K) $\downarrow$ FLOPs (M) $\downarrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$
III-D2 Number of BMSS Blocks
@@
III-D3 Compression Factor in Bottleneck Mechanism
@@ IV Conclusion
This paper presents the first frame-level error annotations in an in-vivo surgical dataset. More than 970000 frames from 48 videos in SAR-RAP50 are annotated frame by frame to provide temporal boundaries of errors. Based on this, we propose SEDMamba, a hierarchical surgical error detection model based on selective SSM with bottleneck and fine-to-coarse temporal fusion, which demonstrates improvements in error detection performance with high computational efficiency. Experimental results show that the proposed method outperforms six state-of-the-art competitions with 1.82-13.06% AUC and 3.8-18.86% AP performance improvements and better qualitative results. Exploiting a very efficient design SEDMamba requires between 11.71-75.79% less model parameters leading to a 40.71-90.93% FLOPs reduction. Future work will focus on detecting error types semantically with multi-label algorithms.

### v2

@@ III-B Evaluation Metrics
Figure 4: Visualization results of the proposed SEDMamba (green line) and the second-best method Mamba (blue line). Each row represents one complete test video. The X-axis is the frame index while the Y-axis is the error probability output by the models. The red translucent background indicates error frames, while the white background indicates normal frames, i.e. ground truth of each frame. Different error types are considered, along with the corresponding prediction results (i.e., error probability). Red and black boxes are used to highlight significant results.
TABLE II: Quantitative results of SEDMamba and other methods for all test data and different error duration ranges. Note that all metrics are calculated on frame level.
Methods All Test Data Short Errors ($<$ 3 seconds) Long Errors ($\geq$ 3 seconds)
AUC (%) $\uparrow$ AP (%) $\uparrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$
LSTM [22] 68.43$\pm$1.21 37.26$\pm$2.65 74.36$\pm$0.90 33.06$\pm$1.27 61.54$\pm$0.62 21.34$\pm$1.56
TeCNO [35] 58.14$\pm$0.33 26.01$\pm$0.32 57.95$\pm$1.90 22.23$\pm$1.41 58.72$\pm$0.90 16.00$\pm$0.20
MS-TCN [36] 58.82$\pm$0.16 29.98$\pm$3.06 55.73$\pm$1.29 29.06$\pm$0.39 61.26$\pm$2.94 14.93 $\pm$1.10
MS-TCN++ [37] 58.57$\pm$0.43 30.35$\pm$1.35 55.34$\pm$1.26 26.48$\pm$1.95 61.44$\pm$0.48 17.99$\pm$1.53
ASFormer [24] 68.76$\pm$0.93 36.79$\pm$2.35 75.07$\pm$1.34 32.70$\pm$2.24 61.72$\pm$2.24 20.14$\pm$0.66
Mamba [27] 69.38$\pm$0.26 41.07$\pm$0.95 75.73$\pm$0.42 35.12$\pm$1.72 62.41$\pm$0.74 25.10$\pm$1.78
Vim [28] 69.06$\pm$1.46 40.63$\pm$2.30 75.24$\pm$0.88 34.64$\pm$1.54 62.02$\pm$2.30 24.67$\pm$1.08
SEDMamba (Ours) 71.20$\pm$0.26 44.87$\pm$1.52 77.92$\pm$0.76 38.91$\pm$0.57 64.01$\pm$0.85 26.71$\pm$0.50
@@ G Params (K) $\downarrow$ FLOPs (M) $\downarrow$ AUC (%) $\uparrow$ AP (%) $\uparrow$
III-C Comparison with the State-of-the-Art
We compare the proposed SEDMamba with state-of-the-art SED method: LSTM [22]; convolutional and Transformer-based video analysis methods: TeCNO [35], MS-TCN [36], MS-TCN++ [37], ASFormer [24]; and selective SSM-based methods: Mamba [27] and Vim [28]. All methods are implemented based on their released code and original literature, and fine-tuned to fit our error detection task to provide a basis for a fair comparison.
Table II shows quantitative results of all methods. For the overall performance (i.e. all test data regardless of error duration), the proposed SEDMamba consistently achieves superior performance with an increase of 1.82-13.06% in AUC and 3.8-18.86% in AP, compared to other methods. Notably, all selective SSM-based methods surpass the Transformer-based method ASFormer, demonstrating the superiority of selective SSM in the SED task. Moreover, SEDMamba achieves gains of 2.19-22.58% in AUC and 3.79-16.68% in AP for errors with short duration ($<$ 3 seconds), and 1.6-5.29% in AUC and 1.61-11.78% in AP for errors with long duration ($>$ 3 seconds).
Fig. 4 illustrates typical error detection results on test videos. [^1] Additional detection results via various error duration are included in the supplementary video. The second-best method Mamba tends to output smoother and more ambiguous prediction curves, which impedes its sensitivity to errors. It also produces lower confidence outputs within error frames. Thanks to the fine-to-coarse temporal information provided by FCTF for selective SSM, SEDMamba is able to identify suddenly occurring errors with short duration (highlighted in red boxes) and consistently produces high confidence within error areas in most cases (highlighted in black boxes). In most normal frames, SEDMamba has comparable or lower error probabilities to Mamba, with only a small portion showing higher error probabilities. Note that both AUC and AP comprehensively consider all classes, i.e. ‘error’ and ‘normal’, indicating SEDMamba has superior performance in both classes.
III-D Computational Efficiency
Given a video spatial embedding sequence $\mathcal{E}\in\mathcal{R}^{L\times D}$, the computational complexity of a single-head self-attention Transformer and a selective SSM in SEDMamba are $O(L^{2}D+LD^{2})$ and $O(LDN)$, respectively. Computationally, self-attention in Transformer is quadratic to sequence length $L$, while selective SSM is linear to $L$. This computational efficiency makes SEDMamba scalable for gigapixel applications with large video lengths. Besides, as shown in Table II and Table III, SEDMamba achieves the highest performance yet maintains the lowest parameters and FLOPs. Specifically, SEDMamba achieves a 91.67% reduction in parameters and a 92.32% reduction in FLOPs compared to the cutting-edge SED method (LSTM), while delivering an AUC improvement of 2.77% and an AP increase of 7.61%. When compared to methods based on 1D and dilated convolution, i.e. TeCNO, MS-TCN and MS-TCN++, our method gets far more performance improvement with a smaller complexity. Furthermore, the parameters and FLOPs of our SEDMamba are only 27.38% and 9.07% of those of the Transformer-based method ASFormer, respectively, yet it achieves a performance gain of 2.44% in AUC and 8.08% in AP. Against other selective SSM-based methods, Mamba and Vim, our SEDMamba obtains 58.23% and 75.79% model parameter reductions, and 51.79% and 70.90% FLOPs reductions, respectively. This is attributed to the bottleneck construction of BMSS blocks, which compresses the spatial dimension at the input and restores it at the output for computational saving, as well as the FCTF therein BMSS, which provides fine-to-coarse temporal features for selective SSM to model long-term dependencies effectively.
III-E Ablation Study
III-E1 Effectiveness of Key Components in BMSS Block
We study the effect of bottleneck mechanism and FCFT in BMSS blocks. As shown in Table IV, whereby removing each component leads to performance degradation. The bottleneck mechanism yields performance gains of +1.64% in AUC and +3.69% in AP, while model parameters and FLOPs are significantly decreased by 82.98% and 78.83%, which guards against overfitting and enhances model efficiency. Besides, such a bottleneck mechanism compresses features on the spatial dimension, which can mitigate the impact of redundant spatial noise on long-term dependency modelling. Introducing FCTF to bottleneck mechanism leads to a slight increase in computational complexity, which is acceptable given the substantial performance gains of +1.58% in AUC and +3.58% in AP. This underscores the importance of providing multi-scale temporal information for selective SSM in the SED task.
III-E2 Number of BMSS Blocks
@@
III-E3 Compression Factor in Bottleneck Mechanism
@@ IV Conclusion
In this paper, we build a novel SED dataset by providing the first-of-its-kind, frame-level error annotations of in-vivo RAS cases. More than 970000 frames from 48 videos in SAR-RAP50 are annotated to provide temporal boundaries of surgical errors. We then propose SEDMamba, a hierarchical SED model based on selective SSM with bottleneck and fine-to-coarse temporal fusion, which yields significant improvements in error detection performance with high computational efficiency. Experimental results show that the proposed method outperforms seven state-of-the-art competitions with 1.82-13.06% AUC and 3.8-18.86% AP performance improvements and better qualitative results. Exploiting a very efficient design SEDMamba requires between 11.71-91.67% less model parameters leading to a 40.71-92.32% FLOPs reduction. Future work will focus on detecting error types semantically with multi-label algorithms.